# Error thresholds of toric codes with transversal logical gates

### v2

@@ IV Toric codes under tCNOT gate with syndrome errors
To model syndrome errors, we assume that each stabilizer measurement has a probability $q$ of being incorrect. In addition, we assume that physical bit-flip errors occur with probability $p$ only between the tCNOT gate and the next syndrome extraction (but not between the syndrome extraction and the tCNOT gate), see Fig. 8. This error model singles out the propagation of syndrome error by the tCNOT gates.
@@
Here, each individual Hamiltonian on the RHS is the one of the R4bIM given in Eq. (II.2), while $t\geq T$ in $H_{3D}^{t\leq T}$ means that this Hamiltonian only contains terms that are defined on the plaquettes at or before $t=T$, and similarly for $H_{3D}^{t>T}$. In fact, the Hamiltonian $H_{3D}(\{\sigma\},\{\tau\}|E^{c},E^{t})$ represents two decoupled R4bIMs with a spin permutation defect plane at $t=T$ that exchanges $\tau\leftrightarrow\tau\sigma$. Redefining $\tau$ to be $\tau\sigma$ at $t>T$ in the partition function further decouples the two flavors of spins there. As a result, they are only coupled via the following set of 5-body interaction terms between $t=T$ and $t=T+1$ (see Fig. 9(a)):
$-J\sum_{l}s^{t}_{l}(T+\frac{1}{2})\sigma_{l}(T)\tau_{l}(T)\tau_{l}(T+1)\tau_{l_{1}}(T+\frac{1}{2})\tau_{l_{2}}(T+\frac{1}{2}).$ (22)
Due to the coupling terms in Eq. (22), we expect the error threshold of the target to be modified near $t=T$. However, it is difficult to obtain a directly analytic statement on the error threshold, and a numerical simulation of the entire stat-mech model in Eq.(IV) is required. Nevertheless, techniques in statistical mechanics like high- and low temperature expansion often shed light towards critical points through its ability of connecting seemingly different stat-mech models [51]. This includes the renowned Kramers-Wannier duality in the 2D Ising model, where the critical point can be computed exactly by comparing the expressions of high- and low temperature expansion.
Following these inspirations, we perform a high temperature expansion of the terms involving the $\{\tau\}$ spins in the Hamiltonian, with the assumption that the $\{\sigma\}$ spins are in the deconfined (decodable) phase when the $\{\tau\}$ spins reach the critical point. In this way, the partition function in Eq. (IV) can be written as
@@
Here, the summation above is over closed surface $\mathfrak{S}$ on the 3D cubic lattice. $\tanh^{|\mathfrak{S}|}(J,K)$ is a shorthand notation that represents the product of $|\mathfrak{S}|$ factors of $\tanh J$ and $\tanh K$. The factor $4$ comes from the fact that each classical spin is shared by two plaquettes, therefore summing over them produces a factor of $2^{2}=4$. Similarly, $\prod_{p\in\mathfrak{S}}(s^{t}_{p},r^{t}_{p})$ represents the product of the random signs of $s^{t}$ and $r^{t}$ in the Hamiltonian terms. $Z_{c}(E^{c})=\sum_{\{\sigma\}}e^{-H_{3D}(\{\sigma\}|E^{c})}$ is the partition function of a single R4bIM, and $\overline{\cdots}$ denote the thermally averaged value with respect to $H_{3D}(\{\sigma\}|E^{c})$. $\mathfrak{S}\cap(t=T)$ denotes the loop at the intersection between the surface $\mathfrak{S}$ and the time slice $t=T$, see Fig. 9(b). Since the $\{\sigma\}$ spins are in the deconfined phase, we expect that the thermal average for a typical $E^{c}$ follows the perimeter-law, so that $\overline{\prod_{l\in\gamma}\sigma_{l}}\sim e^{-A|\gamma|}$ for any loop $\gamma$ with some constant $A>0$. Therefore, when the $\{\sigma\}$ spins are summed over, the $\{\tau\}$ spins are in an effective R4bIM, whose coupling constant $J$ is modified as
@@
for all the terms in Eq. (22) that are located near the plane defect. Intuitively, this effect can be regarded as an increase of local temperature near $t=T$. Hence, with increasing error rates $p$ and $q$, the $\{\tau\}$ spins near $t=T$ will enter a local confined phase before the confinement transition for the rest of the system. From the viewpoint of decoding, this implies that the target code block has a finite but lower error threshold near $t=T$ compared to the rest of the system.
For a rough, conservative estimate for the local error threshold of the target block near $t=T$, we consider the situation where the Wilson loop of the $\{\sigma\}$ spin renormalizes the coupling constant $J$ to $J_{\text{eff}}$ in a considerable region around $t=T$. If we limit ourselves to the case where the syndrome error rate $q$ equal the bit-flip rate $p$, setting $J=K$, the target block reaches threshold when $J_{\text{eff}}(p)$ reaches its critical value:
@@
This implies that the reduction of error threshold from that of memory with imperfect syndromes, $p^{*}=0.033$ will be moderate as long as the loop tension $A$ is small. The numerical study of R4bIM in Ref. [10] found that $A$ is of order $0.01$ when $p$ is close but smaller than $p^{*}$. Therefore, we estimate that the solution to Eq. (25) by setting $A=0.01$, which yields $p_{\text{c}}^{t}=0.028$. Hence, even with conservative assumptions, the decrease in the error threshold of the target block due to the spread of syndrome error is not significant. We leave the direct numerical simulation of the local error threshold for future work.
V Conclusion and outlook
@@
Figure 10: (a) A noisy transversal Clifford logical circuit between two code blocks $A$ and $B$ consisting of two tCNOT gates and a (fold-)transversal $S$ gate. Ever pair of transversal gates are separated by $\Delta t$ rounds of syndrome error extractions and generic physical error channels $\mathcal{N}$. (b) Sketch of the impact of transversal gates on the stat-mech model that is mapped from the circuit in (a). The $X/Z$ spacetime detectors of the two code blocks $A$ and $B$, denoted by $\{d^{A/B}_{X/Z}\}$, give rise to 4 different set of classical spins in total (see Appendix C for details), which are coupled at the plane defects induced by the logical operations. The propagation of Pauli errors by the transversal gates are marked by red arrows. The impact region $|t-T|<\lambda$ due to increased effective local temperature are marked by darker colors for each plot.
@@
In summary, we studied the error threshold of the tCNOT gate in toric codes under bit-flip noise using the stat-mech mapping approach. We considered two error models, the persistent bit-flip error and bit-flip plus syndrome errors. These error models quantify the reduction of threshold compared to quantum memory due to two distinct error spreading mechanisms by the tCONT gate: the spread of physical bit-flip error and syndrome error. In case of persistent bit-flip error with perfect syndromes, we mapped the decoding problem to a 2D random AT model. The error thresholds of control and target code blocks in this case were $p_{\text{c}}^{c}=0.099$ and $p_{\text{c}}^{t}=0.080$, respectively. These values were rigorously established via Monte-Carlo simulation and FSS. The case with syndrome error was mapped to a plane defect in 3D R4bIM, and the decrease in the error threshold of the target block, $p_{\text{c}}^{t}=0.028$, was conservatively estimated leveraging previously reported numerical results [10]. Comparing these thresholds to the memory thresholds in respective cases, our results demonstrate that both mechanisms of error spreading only cause a minor, non-devastating decrease in the decoding threshold compared to the memory threshold.
Throughout this work, we specifically consider toric code logical circuits with a single tCNOT gate. However, our study can be straightforwardly generalized to map a general transversal Clifford logical circuit of toric codes to a stat-mech model, where each consecutive logical operation is separated by $\Delta t$ rounds of syndrome extractions, see Fig. 10(a). In Appendix C, we carry out detailed stat-mech mapping of fold-transversal Hadamard and $S$ gate for toric code. Combined with the tCNOT gate discussed in the main text, we are able to derive stat-mech model of the full transversal Clifford logical circuit. Crucially, each transversal Clifford gate operation affects a specific time slice in the 3D stat-mech model. Hence, the effect of each gate at $t=T$ will be local in the time direction inside a time window $|t-T|\leq\lambda$, where $\lambda\sim O(1)$, as shown pictorially in Fig. 10(b). For a given circuit, one should be able to simulate the corresponding models based on the mappings that we have provided in this paper. Therefore, our estimation of error threshold can be readily generalized to estimate the threshold of the entire logical circuit.
Our stat-mech approach can even be generalized to arbitrary Calderbank-Steane-Shor (CSS) codes that possess transversal logical gates. We refer interested readers to Appendix D for more details.
Interestingly, our success in mapping transversal logical circuits to stat-mech models signals a potential connection between the error threshold of the logical circuit and topological phase transitions. Namely, the entire noisy logical circuit within and beyond the error threshold can be regarded as two distinct phases of matter. The determination of error threshold of logical circuit can then be formulated as a phase-recognition task suitable for learning agents [52].
@@ Appendix A Coupling constants in the Hamiltonian of the random Ashkin-Teller model
In this appendix, we derive the coupling constants $K_{2}$ and $K_{4}$ in the random AT model in Eq. (17). To this end, we first rewrite the probability distribution of the random link distribution in Eq. (III.3) in an exponential ansatz:
@@ Appendix B Derivation of the stat-mech model for tCNOT with syndrome errors
To parameterize the error $E$ with syndrome errors, we use a binary number $r^{c,t}_{p}(t)=1/-1$ to denote that the syndrome extraction of the $Z$ stabilizer at the plaquette $p$ in the control or target block is correct/incorrect, respectively. $t\in\mathbb{Z}$ in the bracket is the time step at which the syndrome extraction happens. The probability distribution is given by $p_{r}(r^{c,t}_{p}(t))=\frac{1+(1-2q)r^{c,t}_{p}(t)}{2}$. The bit-flip errors are now represented by $s_{l}^{c,t}(t+\frac{1}{2})=\pm 1$ with probability $p_{s}(s_{l}^{c,t}(t+\frac{1}{2}))=\frac{1+(1-2p)s_{l}^{c,t}(t+\frac{1}{2})}{2}$. Without the tCNOT gate, the values of the spacetime detectors are given by
@@
Figure 11: The stat-mech mapping of the toric code with generic Pauli and syndrome errors. Here the physical qubits of the toric code are marked by black circles, the shaded and white squares denote weight-4 $Z$ and $X$ stabilizers, respectively. The two flavors of Ising spins, $\{\sigma^{x}\}$ and $\{\sigma^{z}\}$, each live on links of a cubic lattice. As such, at $t\in\mathbb{Z}+\frac{1}{2}$, the Ising spins live on two 2D square lattices that are dual to each other, and at $t\in\mathbb{Z}$ they overlap at the same spacial locations.
@@ Appendix C Stat-mech mapping for transversal Clifford gates and generic Pauli error channels
where $i$ is the index of qubits. In this way, the bit-flip channel with probability $p$ corresponds to $(p_{x},p_{y},p_{z})=(p,0,0)$ and the depolarizing channel with probability $p$ corresponds to $(p_{x},p_{y},p_{z})=(p/3,p/3,p/3)$. To facilitate the stat-mech mapping, we can represent the absence/occurrence of single-qubit Pauli error as two correlated random variables $s^{x}_{i}=\pm 1$ and $s^{z}_{i}=\pm 1$, which follows the probability distribution
@@
To model syndrome extraction errors, we consider an error model in which every $X$ and $Z$ stabilizer has probabilities $q_{x}$ and $q_{z}$ of being the opposite of its actual value, respectively. The absence/presence of such errors can be parameterized by parities $r^{x}=\pm 1$ and $r^{z}=\pm 1$. They follow the probability distributions
@@
$s^{x}_{l}(t-\frac{1}{2})\to s^{x}_{l}(t-\frac{1}{2})\sigma^{x}_{l}(t-1)\sigma^{x}_{l}(t)\sigma^{x}_{l_{1}}(t-\frac{1}{2})\sigma^{x}_{l_{2}}(t-\frac{1}{2}),\ s^{z}_{l}(t-\frac{1}{2})\to s^{z}_{l}(t-\frac{1}{2})\sigma^{z}_{l}(t-1)\sigma^{z}_{l}(t)\sigma^{z}_{l_{1}}(t-\frac{1}{2})\sigma^{z}_{l_{2}}(t-\frac{1}{2}),$
@@
Using the standard steps of stat-mech mapping outlined in Sec. II, the toric code quantum memory under the generic Pauli errors can be mapped to a stat-mech model described by the following partition function

### v3

@@ 4 Toric codes under tCNOT gate with syndrome errors
To model syndrome errors, we assume that each stabilizer measurement has a probability $q$ of being incorrect. In addition, we assume that physical bit-flip errors occur with probability $p$ only between the tCNOT gate and the next syndrome extraction (but not between the syndrome extraction and the tCNOT gate), see Fig. 8. This error model focuses on the propagation of syndrome errors by the tCNOT gates.
@@
Here, each individual Hamiltonian on the RHS is the one of the R4bIM given in Eq. (2.2), while $t\geq T$ in $H_{3D}^{t\leq T}$ means that this Hamiltonian only contains terms that are defined on the plaquettes at or before $t=T$, and similarly for $H_{3D}^{t>T}$. In fact, the Hamiltonian $H_{3D}(\{\sigma\},\{\tau\}|E^{c},E^{t})$ represents two decoupled R4bIMs with a spin permutation defect plane at $t=T$ that exchanges $\tau\leftrightarrow\tau\sigma$. Redefining $\tau$ to be $\tau\sigma$ at $t>T$ in the partition function further decouples the two flavors of spins there. As a result, they are only coupled via the following set of 5-body interaction terms between $t=T$ and $t=T+1$ (see Fig. 9(a)):
$-J\sum_{l}$ $s^{t}_{l}(T+\frac{1}{2})\sigma_{l}(T)\times$
$\tau_{l}(T)\tau_{l}(T+1)\tau_{l_{1}}(T+\frac{1}{2})\tau_{l_{2}}(T+\frac{1}{2}).$ (22)
Due to the coupling terms in Eq. (4), we expect the error threshold of the target to be modified near $t=T$. However, obtaining a direct analytic statement of the error threshold is difficult, and a numerical simulation of the entire stat-mech model in Eq. (4) is required. Nevertheless, techniques in statistical mechanics, such as high- and low-temperature expansions, often shed light on critical points by connecting seemingly different stat-mech models [51]. This includes the renowned Kramers-Wannier duality in the 2D Ising model, where the critical point can be computed exactly by comparing the high- and low-temperature expansions.
Following these inspirations, we perform a high-temperature expansion of the terms involving the $\{\tau\}$ spins in the Hamiltonian, with the assumption that the $\{\sigma\}$ spins are in the deconfined (decodable) phase when the $\{\tau\}$ spins reach the critical point. In this way, the partition function in Eq. (4) can be written as
@@
Here, the summation above is over the closed surface $\mathfrak{S}$ on the 3D cubic lattice. $\tanh^{|\mathfrak{S}|}(J,K)$ is a shorthand notation that represents the product of $|\mathfrak{S}|$ factors of $\tanh J$ and $\tanh K$. The factor $4$ arises because each classical spin is shared by two plaquettes; summing over them yields a factor of $2^{2}=4$. Similarly, $\prod_{p\in\mathfrak{S}}(s^{t}_{p},r^{t}_{p})$ represents the product of the random signs of $s^{t}$ and $r^{t}$ in the Hamiltonian terms. $Z_{c}(E^{c})=\sum_{\{\sigma\}}e^{-H_{3D}(\{\sigma\}|E^{c})}$ is the partition function of a single R4bIM, and $\overline{\cdots}$ denote the thermally averaged value with respect to $H_{3D}(\{\sigma\}|E^{c})$. $\mathfrak{S}\cap(t=T)$ denotes the loop at the intersection between the surface $\mathfrak{S}$ and the time slice $t=T$, see Fig. 9(b). Since the $\{\sigma\}$ spins are in the deconfined phase, we expect that the thermal average for a typical $E^{c}$ follows the perimeter-law, so that $\overline{\prod_{l\in\gamma}\sigma_{l}}\sim e^{-A|\gamma|}$ for any loop $\gamma$ with some constant $A>0$. Therefore, when the $\{\sigma\}$ spins are summed over, the $\{\tau\}$ spins are in an effective R4bIM, whose coupling constant $J$ is modified as
@@
for all the terms in Eq. (4) that are located near the plane defect. Intuitively, this effect can be regarded as an increase in local temperature near $t=T$. Hence, with increasing error rates $p$ and $q$, the $\{\tau\}$ spins near $t=T$ will enter a local confined phase before the confinement transition for the rest of the system. From the decoding perspective, this implies that the target code block has a finite but lower error threshold near $t=T$ than the rest of the system.
For a rough, conservative estimate for the local error threshold of the target block near $t=T$, we consider the situation where the Wilson loop of the $\{\sigma\}$ spin renormalizes the coupling constant $J$ to $J_{\text{eff}}$ in a considerable region around $t=T$. If we limit ourselves to the case where the syndrome error rate $q$ equals the bit-flip rate $p$, setting $J=K$, the target block reaches threshold when $J_{\text{eff}}(p)$ reaches its critical value:
@@
This implies that the reduction of error threshold from that of memory with imperfect syndromes, $p^{*}=0.033$ will be moderate as long as the loop tension $A$ is small. The numerical study of R4bIM in Ref. [10] found that $A$ is of order $0.01$ when $p$ is close but smaller than $p^{*}$. Therefore, we estimate the solution to Eq. (25) by setting $A=0.01$, which yields $p_{\text{c}}^{t}=0.028$. Hence, even under conservative assumptions, the decrease in the target block’s error threshold due to the spread of syndrome errors is not significant. We leave the direct numerical simulation of the local error threshold for future work.
5 Conclusion and outlook
@@
Figure 10: (a) A noisy transversal Clifford logical circuit between two code blocks $A$ and $B$ consisting of two tCNOT gates and a (fold-)transversal $S$ gate. Every pair of transversal gates is separated by $\Delta t$ rounds of syndrome error extractions and generic physical error channels $\mathcal{N}$. (b) Sketch of the impact of transversal gates on the stat-mech model that is mapped from the circuit in (a). The $X/Z$ spacetime detectors of the two code blocks $A$ and $B$, denoted by $\{d^{A/B}_{X/Z}\}$, give rise to 4 different sets of classical spins in total (see Appendix C for details), which are coupled at the plane defects induced by the logical operations. The propagation of Pauli errors by the transversal gates is marked by red arrows. The impact region $|t-T|<\lambda$ due to increased effective local temperature is marked by darker colors for each plot.
In this work, we have generalized the statistical-mechanical mapping, a foundational tool for establishing rigorous, decoder-independent error thresholds for quantum memories, to logical circuits with transversal gates. Our work provides the first extension of stat-mech mapping to circuits that perform logical operations. The key insight underlying this generalization is that each transversal gate introduces a permutation defect of the classical spins at the time slice where the gate is applied, modifying the stat-mech model only locally in time.
@@
We demonstrated this framework by deriving stat-mech models for two toric code blocks undergoing a tCNOT gate under two complementary error models. For persistent bit-flip errors with perfect syndromes, the decoding problem maps to a 2D random AT model, and Monte Carlo simulations with FSS yield error thresholds of $p_{\text{c}}^{c}=0.099$ and $p_{\text{c}}^{t}=0.080$ for the control and target code blocks, respectively. For bit-flip errors coexisting with syndrome errors, the circuit maps to a 3D R4bIM with a plane defect, and we conservatively estimate the target block threshold at $p_{\text{c}}^{t}\geq 0.028$. Comparing with the corresponding memory thresholds ($p_{\text{th}}=0.109$ and $p^{*}=0.033$), these results establish that the error spreading by the tCNOT gate causes a moderate, non-devastating decrease in the decoding threshold—a reassuring conclusion for the viability of transversal-gate-based fault-tolerant computation.
An essential feature of our framework is its systematic generalizability. In Appendix C, we derive the stat-mech models for the fold-transversal Hadamard and $S$ gates of the toric code under generic single-qubit Pauli noise, showing that the transversal Hadamard gate induces a flavor-exchange defect ($\sigma^{x}\leftrightarrow\sigma^{z}$), while the transversal $S$ gate introduces a coupling between the $X$ and $Z$ error sectors identical in structure to the tCNOT plane defect. Together with the tCNOT gate, these results provide a complete stat-mech mapping for any transversal Clifford logical circuit of the toric code, where each consecutive gate is separated by $\Delta t$ rounds of syndrome extractions, as illustrated in Fig. 10(a). Crucially, each gate affects the stat-mech model only within a time window $|t-T|\leq\lambda$ with $\lambda\sim O(1)$, as sketched in Fig. 10(b). This locality ensures that the threshold of the full circuit can be estimated by simulating the stat-mech model with all the plane defects corresponding to the gates in the circuit.
In Appendix D, we further extend the framework to arbitrary CSS codes that support transversal Clifford gates, using a binary matrix formalism. We prove that: (i) every transversal gate modifies the spacetime parity check matrix via a local modification that changes the kernel of the check matrix; (ii) the resulting stat-mech model differs from the memory model only at the time of the gate; and (iii) the effects of multiple transversal gates compose multiplicatively (Eq. (D.1)). These results establish that the stat-mech mapping is, in principle, applicable to any transversal Clifford logical circuit built from CSS codes, providing a systematic route to rigorous threshold analysis of fault-tolerant quantum computation.
Looking ahead, our work opens several exciting directions. The stat-mech models derived here can be simulated directly for specific circuits of interest, thereby enabling threshold benchmarks for near-term fault-tolerant architectures. Furthermore, the connection between error thresholds and phase transitions in the stat-mech model suggests a potential link to topological phases of matter: the noisy logical circuit within and beyond the error threshold may be viewed as two distinct phases, and the determination of the threshold can be formulated as a phase-recognition task amenable to machine learning approaches [52]. Finally, extending the stat-mech mapping beyond Clifford circuits, potentially by incorporating magic-state injection or non-Clifford transversal gates, remains an important open problem.
@@ Appendix A Coupling constants in the Hamiltonian of the random Ashkin-Teller model
In this appendix, we derive the coupling constants $K_{2}$ and $K_{4}$ in the random AT model in Eq. (17). To this end, we first rewrite the probability distribution of the random link distribution in Eq. (3.3) in an exponential ansatz:
@@ Appendix B Derivation of the stat-mech model for tCNOT with syndrome errors
To parameterize the error $E$ with syndrome errors, we use a binary number $r^{c,t}_{p}(t)=1/-1$ to denote that the syndrome extraction of the $Z$ stabilizer at the plaquette $p$ in the control or target block is correct/incorrect, respectively. $t\in\mathbb{Z}$ in the brackets denotes the time step at which syndrome extraction occurs. The probability distribution is given by $p_{r}(r^{c,t}_{p}(t))=\frac{1+(1-2q)r^{c,t}_{p}(t)}{2}$. The bit-flip errors are now represented by $s_{l}^{c,t}(t+\frac{1}{2})=\pm 1$ with probability $p_{s}(s_{l}^{c,t}(t+\frac{1}{2}))=\frac{1+(1-2p)s_{l}^{c,t}(t+\frac{1}{2})}{2}$. Without the tCNOT gate, the values of the spacetime detectors are given by
@@
Figure 11: The stat-mech mapping of the toric code with generic Pauli and syndrome errors. Here, the physical qubits of the toric code are marked by black circles, the shaded and white squares denote weight-4 $Z$ and $X$ stabilizers, respectively. The two flavors of Ising spins, $\{\sigma^{x}\}$ and $\{\sigma^{z}\}$, each live on links of a cubic lattice. As such, at $t\in\mathbb{Z}+\frac{1}{2}$, the Ising spins live on two 2D square lattices that are dual to each other, and at $t\in\mathbb{Z}$ they overlap at the same spatial locations.
@@ Appendix C Stat-mech mapping for transversal Clifford gates and generic Pauli error channels
where $i$ is the index of qubits. In this way, the bit-flip channel with probability $p$ corresponds to $(p_{x},p_{y},p_{z})=(p,0,0)$ and the depolarizing channel with probability $p$ corresponds to $(p_{x},p_{y},p_{z})=(p/3,p/3,p/3)$. To facilitate the stat-mech mapping, we can represent the absence/occurrence of a single-qubit Pauli error as two correlated random variables $s^{x}_{i}=\pm 1$ and $s^{z}_{i}=\pm 1$, which follows the probability distribution
@@
To model syndrome-extraction errors, we consider an error model in which each $X$ and $Z$ stabilizer has probabilities $q_{x}$ and $q_{z}$ of being the opposite of its actual value, respectively. The absence/presence of such errors can be parameterized by parities $r^{x}=\pm 1$ and $r^{z}=\pm 1$. They follow the probability distributions
@@
$s^{x}_{l}(t-\frac{1}{2})\to s^{x}_{l}(t-\frac{1}{2})\sigma^{x}_{l}(t-1)\sigma^{x}_{l}(t)\sigma^{x}_{l_{1}}(t-\frac{1}{2})\sigma^{x}_{l_{2}}(t-\frac{1}{2}),$
$s^{z}_{l}(t-\frac{1}{2})\to s^{z}_{l}(t-\frac{1}{2})\sigma^{z}_{l}(t-1)\sigma^{z}_{l}(t)\sigma^{z}_{l_{1}}(t-\frac{1}{2})\sigma^{z}_{l_{2}}(t-\frac{1}{2}),$
@@
Using the standard steps of stat-mech mapping outlined in Sec. 2, the toric code quantum memory under the generic Pauli errors can be mapped to a stat-mech model described by the following partition function